# Modularity aided consistent attributed graph clustering via coarsening

### v1

@@ 1 Introduction
Cut-based methods (Wei & Cheng, 1989; Shi & Malik, 2000; Bianchi et al., 2020), aiming to minimize the number of edges (or similar metric) in a cut, may fall short in capturing the true community structure if the cut’s edge count doesn’t significantly deviate from random graph expectations. This approach originated from the Fiedler vector, which yields a graph cut with a minimal number of edges over all possible cuts (Fiedler, 1973; Newman, 2006a). Similarity-based techniques, reliant on pairwise node similarities for clustering leverage metrics of similarity to group nodes with shared characteristics. These can be computationally intensive and susceptible to noise, potentially yielding suboptimal results, especially in sparse or noisy data scenarios. Modularity-based methods rely on a statistical approach, measuring the disparity in edge density between a graph and a random graph with the same degree sequence. On the other hand, these methods exhibit a resolution limit (Fortunato & Barthélemy, 2007), as the emphasis on maximizing modularity may inadvertently lead to the neglect of smaller community structures within the graph. Each clustering approach, whether cut-based, similarity-based, or modularity-based, bears its own set of limitations that necessitates careful consideration based on the characteristics of the underlying graph data.
Moreover, graph reduction techniques such as coarsening, summarization, or condensation can also be utilized to facilitate the clustering task (Dhillon et al., 2007; Kumar et al., 2023; Loukas & Vandergheynst, 2018; Loukas, 2019). In the context of graph coarsening, the objective is to learn a reduced graph by merging similar nodes into supernodes. While this coarsening process is traditionally employed for graph reduction, it can be extended to clustering by reducing the original graph such that each class corresponds to a supernode. However, relying solely on the loss function of coarsening may lead to suboptimal performance, particularly when reducing the graph size to the number of classes results in a considerable reduction and may lead to significant loss of information.
In this study, we introduce an optimization-based framework designed to enhance the clustering task by leveraging both the adjacency and the node features of the graph. Our proposed framework strategically incorporates coarsening and modularity maximization, refining partitioning outcomes and bolstering the overall effectiveness of the clustering process. The algorithm minimizes a nuanced loss function, encompassing a $\log\det$ term, smoothness, and modularity components, to ensure efficient clustering. Formulated as a multi-block non-convex optimization problem, our approach is adeptly addressed through a block majorization-minimization technique, wherein variables are updated individually while keeping the remainder fixed. The resulting algorithm demonstrates convergence, showcasing its efficacy in efficiently solving the proposed optimization problem.
@@
We introduce two additional algorithms, Q-VGAE and Q-GMM-VGAE, incorporating variational graph auto-encoders to further enhance clustering accuracy. Through comprehensive experiments, we demonstrate the effectiveness of our proposed algorithms, clearly surpassing the performance of current state-of-the-art methods on benchmark datasets. Our innovative approach, showcases notable improvements in clustering performance, thereby solidifying the robustness and superiority of our proposed methods in handling real-world benchmark datasets.
@@
Notations. Let $G=\{V,E,A,X\}$ be a graph with node set $V=\{v_{1},v_{2},...,v_{p}\}\ (|V|=p)$, edge set $E\subset V\times V\}(|E|=e)$, weight (adjacency) matrix $A$ and node feature matrix $X\in\mathbb{R}^{p\times n}$. Also, let $\textbf{d}=A.\textbf{1}\in\mathbb{Z}_{+}^{p}$ be the degree vector, where 1 is the vector pf size $p$ having all entries 1. Then, the graph Laplacian is $\Theta=\textrm{diag}(\textbf{d})-A$ and the set of all valid Laplacian matrices is defined as: $S_{\Theta}=\{\Theta\in\mathbb{R}^{p\times p}|\Theta_{ij}=\Theta{ji}\leq 0\ \text{for}\ i\neq j,\Theta_{ii}=\sum_{j=1}^{p}\Theta_{ij}\}$.
@@ 2 Related Works
Graph Clustering via Coarsening. Graph coarsening can be extended to graph clustering by reducing the size of the coarsened graph to the number of classes, i.e., $k=l$. However, in most graphs, the number of classes is very small, and reducing the graph to this extent may lead to poor clustering quality. For example, DiffPool (Ying et al., 2018) learns soft cluster assignments at each layer of the GNN and two extra losses, entropy to penalize the soft assignments and a link prediction based loss. Next, SAGPool (Lee et al., 2019) calculates attention scores and node embeddings to determine the nodes that need to be preserved or removed. Top-k (Gao & Ji, 2019) also works by sparsifying the graph with the learned weights. MinCutPool (Bianchi et al., 2020) formulates a differentiable relaxation of spectral clustering via pooling. However, (Tsitsulin et al., 2023) show that MinCutPool’s orthogonal regularization dominates over the clustering objective and the objective is not optimized. Some disadvantages of these methods are stability and computational complexity in the case of SAGPool and DiffPool and convergence in MinCutPool. To address these challenges, we need a better loss function to perform the clustering task effectively.
@@
Deep Graph Clustering. Previous literature can be classified based on contrastive and non-contrastive methods. On the non-contrastive side, Pan et al. (2018) proposed ARGA and ARVGA, enforcing the latent representations to align to a prior using adversarial learning. By utilizing an attention-based graph encoder and a clustering alignment loss, Wang et al. (2019) propose DAEGC. Liu et al. (2022) design the DCRN model to alleviate representation collapse by a propagation regularization term minimizing the JSD between the latent and its product with normalized $A$. Contrastive methods include AGE (Cui et al., 2020) which builds a training set by adaptively selecting node pairs that are highly similar or dissimilar after filtering out high-frequency noises using Laplacian smoothing. Zhao et al. (2021) propose GDCL to correct the sampling bias by choosing negative samples based on the clustering label.
@@
VGAEs (Kipf & Welling, 2016a) are an increasingly popular class of GNNs that leverage variational inferenceKingma & Welling (2014) for learning latent graph representations in unsupervised settings. They reconstruct the adjacency matrix after passing the graph through an encoder-decoder architecture. Many attempts have been made to use VGAEs with k-means on latent embeddings, but it has been unsuitable for clustering. This is primarily because embedded manifolds obtained from VGAEs are curved and must be flattened before any clustering algorithms using Euclidean distance are applied. Refer to Appendix Section L for a detailed explanation. VGAEs only use a single Gaussian prior for the latent space, whereas clustering requires the integration of meta-priors. Additionally, the inner-product decoder fails to capture locality and cluster information in the formed edges. Several clustering-oriented variants of VGAEs (Mrabah et al., 2022; Hui et al., 2020) have been developed that overcome most of these challenges. GMM-VGAE (Hui et al., 2020) partitions the latent space using a Gaussian Mixture Model and assigns a separate prior for each cluster to better model complex data distributions. Despite the improvement in performance, it’s inner-product decoder cannot capture locality information.
@@
Modularity Maximization. Various heuristic algorithms have been established that solve the NP-hard problem of modularity maximization including sampling, simulated annealing (Guimerà & Nunes Amaral, 2005; Guimerà & Amaral, 2005), and greedy algorithms (Louvain/Leiden) (Newman, 2004; Blondel et al., 2008a). These algorithms require intensive compute and don’t use node features. Modularity maximization using GNNs has also garnered attention recently. DMoN (Tsitsulin et al., 2023) optimizes only for modularity (with a collapse regularization to prevent the trivial solution) but offers no theoretical guarantees about convergence. Modularity-Aware GAEs and VGAEs (Salha-Galvan et al., 2022) use a prior membership matrix using Louvain algorithm and optimize for modularity using RBF kernel as a same-community assignment proxy. DGCLUSTER (Bhowmick et al., 2023) is a semi-supervised method that makes use of either a subset of labels or pairwise memberships as part of their "Auxiliary Information" and a modularity objective.
@@ 3.1 Graph Coarsening
Graph coarsening is a graph dimensionality reduction technique used in large-scale machine learning to construct a smaller or coarsened graph $G_{c}$ from the original graph $G=\{V,E,A,X\}$ while preserving properties of the original graph $G$. Given an original graph $G(V,E,A,X)$, graph coarsening aims to learn a mapping matrix $C\in\mathbb{R}_{+}^{p\times k}$, where $p$ is the number of nodes in the original graph and $k$ is the number of nodes in the coarsened graph. Each non-zero entry of the mapping matrix $C$, i.e., $C_{ij}$, indicates that the $i$-th node of the original graph is mapped to the $j$-th supernode. Moreover, for a balanced mapping, the mapping matrix $C$ belongs to the following set Kumar et al. (2023); Loukas & Vandergheynst (2018); Loukas (2019):
@@
For $C\in\mathcal{C}$, the relationship between the original graph Laplacian $\Theta$, the coarsened graph Laplacian matrix $\Theta_{c}$, and the mapping matrix $C$ is given by $\Theta_{c}=C^{T}\Theta C$.
@@ 3.2 Spectral Modularity Maximization
Spectral Clustering is the most direct approach to graph clustering, where we minimize the volume of inter-cluster edges (i.e., the total number of edges between clusters, not within them). Modularity, introduced in (Newman, 2006b), is the difference between the number of edges within a cluster $\mathbf{C_{i}}$ and the expected number of such edges in a random graph with an identical degree sequence. It is mathematically defined as:
@@
$B$ $=A-\frac{\textbf{dd}^{T}}{2e},\qquad\mathcal{Q}=\frac{1}{2e}Tr(C^{T}BC)$ (3)
Similar to the graph Laplacian matrix $\mathbf{\Theta}$, $\mathbf{B}$ is symmetric and is defined such that its row-sums and column-sums are zero, thereby making $\mathbf{1}$ one of its eigenvectors and 00 the corresponding eigenvalue. These spectral properties of the modularity matrix are also observed in the Laplacian, as noted in Newman (2006b), which is a crucial element in spectral clustering. Modularity is maximized when $u_{1}^{T}s$ is maximized, where $u$ are the eigenvectors of $B$ and $s$ is the community assignment vector, i.e., placing the majority of the summation in $Q$ on the first (and largest) eigenvalue of $B$. Moreover, modularity is closely associated with community detection. These special spectral properties make $\mathbf{B}$ an ideal choice for graph clustering. While modularity maximization has been extensively studied, heuristic algorithms for it are computationally intensive (such as the Newman-Girvan algorithm with $\mathcal{O}(n^{3})$ time complexity). The Louvain/Leiden algorithms improve on this.
@@ 3.3 Motivation and Proposed Problem Formulation
Current graph clustering methods often struggle to accurately capture both intra-cluster and inter-cluster relationships, achieve computational efficiency, and identify smaller communities, thereby limiting their effectiveness. These methods face significant challenges including instability, high computational complexity, and the lack of convergence guarantees, especially when employing coarsening techniques. Modularity maximization, despite extensive study, relies on computationally intensive heuristics and lacks theoretical convergence guarantees. To address these issues, we propose an optimization-driven framework that strategically integrates coarsening and modularity. By incorporating both adjacency and node features, our approach aims to robustly capture intra-cluster and inter-cluster dynamics. This framework ensures stability, guarantees convergence, and consistently delivers superior clustering results with enhanced computational efficiency compared to existing methods. Given original graph $G(V,E,A,X)$, the proposed optimization formulation is:
@@
subject $\text{subject to}\ C$ $\in\mathcal{S_{C}}=\{C\in\mathbb{R}^{p\times l}|C\geq 0,\ \left\|[C^{T}]_{i}\right\|_{2}^{2}\leq 1\}\;\ \forall\ i=1,2\ldots p$ (4)
Here, $C\in\mathbb{R}_{+}^{p\times l}$ represents the clustering matrix to be learned, where each non-zero entry $C_{ij}$ indicates that the $i$-th node is assigned to the $j$-th cluster. The term $f(C,X,\Theta)$ denotes the graph coarsening objective, managing inter-cluster edges. The function $g(C,A)$ represents the modularity objective, improving clustering performance. Additionally, $h(C,\Theta)$ is a regularization term enforcing desirable properties in the clustering matrix, as defined in Equation 1. The overarching goal of the optimization problem 3.3 is to choose $f(C,X,\Theta)$, $g(C,A)$, and $h(C,\Theta)$ such that nodes are optimally clustered and inter-cluster connectivity is maintained. In the next section, we will develop the clustering algorithm, leveraging the coarsening objective and modularity to enhance clustering performance.
@@ 4 Proposed Method
Given a graph $G=\{V,E,A,X\}$, considering $f(C,X_{C},\Theta)=\text{tr}(X_{C}^{T}C^{T}\Theta CX_{C})$, $g(C,A)=-\frac{\beta}{2e}\textrm{tr}(C^{T}BC)$, and $h(C,\Theta)=-\gamma\log\det(C^{T}\Theta C+J)$, to obtain clustering matrix $C$ we formulate the following optimization problem :
@@
subject $\text{subject to}\ C\in\mathcal{S_{C}}$ =$=\{C\in\mathbb{R}^{p\times l}|\left\|C^{T}_{i}\right\|_{2}^{2}\leq 1\}\;\forall i,X=CX_{C}\;\;\text{where, }J=\frac{1}{k}\textbf{1}_{k\times k}$ (5)
The term $\textrm{tr}(X_{C}^{T}C^{T}\Theta CX_{C})$ represents the smoothness or Dirichlet energy of the coarsened graph while $C^{T}\Theta C$ is the Laplacian matrix of the coarsened graph. Minimzing smoothness ensures that the clusters or supernodes with similar features are linked with stronger weights. The term $\textrm{tr}(C^{T}BC)$ corresponds to the graph’s modularity, enhancing the quality of the clusters formed. The term $-\log\det(C^{T}\Theta C+J)$ is crucial for maintaining inter-cluster edges in the coarsened graph. For a connected graph matrix with $l$ super-nodes or clusters, the rank of $C^{T}\Theta C$ is $l-1$. Adding $J$ to $C^{T}\Theta C$ makes $C^{T}\Theta C+J$ a full-rank matrix without altering the row and column space of $C^{T}\Theta C$ (Kumar et al., 2020).
Problem 4 is a multi-block non-convex optimization problem solved using the Block Successive Upper-bound Minimization (BSUM) framework. All terms except modularity are convex in nature, which we prove in Appendix A. We iteratively update $C$ and $X_{C}$ alternately while keeping the other constant. This process continues until convergence or the stopping criteria are met. Since the constraint $X=CX_{C}$ is hard and difficult to enforce, we relax it by adding the term $\frac{\alpha}{2}\|X-CX_{C}\|_{F}^{2}$. This term ensures that each node is assigned to a cluster, leaving no node unassigned. Additionally, when needed, we can add an optional sparsity regularization term $\lambda\|C^{T}\|_{1,2}^{2}$, which can be seen as ensuring that each node is assigned to exactly one cluster, avoiding any overlap in node assignments across clusters.
@@
By using the first-order Taylor series approximation, a majorised function for $f(C)$ at $C^{(t)}$ can be obtained as:
@@
A graph $G(V,E)$ is generated by a DC-SBM with $p$ nodes which belong to $k$ classes. Following the setup in Zhao et al. (2012), each node $v_{i}$ is associated with a label-degree pair $(y_{i},t_{i})$ drawn from a discrete joint distribution $\Pi_{k\times m}$. Here, $t_{i}$ takes values $0\leq x_{1}\leq\dots\leq x_{m}$ with $\mathbb{E}[t_{i}]=1$, and $y_{i}\in[k]$.
@@
Next, we define $O(e)\in\mathbb{Z}_{\geq 0}^{k\times k}$ as the inter-cluster edge count matrix for a given cluster assignment $e\in[k]^{p}$. Here, $O_{ql}(e)=\sum_{ij}A_{ij}\mathbbm{1}\{e_{i}=q,e_{j}=l\}$ denotes the number of edges between clusters $q$ and $l$. Additionally, we represent the class frequency distribution as $\pi\in[0,1]^{k}$, where $\pi_{i}$ indicates the fraction of nodes assigned to cluster $i$.
@@
(Theorem 4.1 from Zhao et al. (2012)) For any $\mathcal{L}(e)=F(\frac{O(e)}{\mu_{p}},f(e))$, if $\pi,P,F$ is uniquely maximized by $S=\mathbb{D}$ and satisfies the regularity conditions, then $\mathcal{L}$ is strongly consistent under DC-SBMs $\frac{\lambda_{p}}{logp}\rightarrow\infty$ and weakly consistent if $\lambda_{p}\rightarrow\infty$:
@@
It is important to note that the paper focuses on maximizing an objective function. Therefore, we consider the negative of our loss function, $-\mathcal{L}_{MAGC}$. Additionally, we only consider the objective function defined in Equation equation 4, as regularizers can be adjusted for different downstream tasks.
@@
We also demonstrate experimental consistency of the optimization objective equation 4 in Appendix I, resulting in complete recovery of the labels. This consistency extends to the objective with additional regularization terms, as detailed therein.

### v2

@@ 1 Introduction
Cut-based methods (Wei & Cheng, 1989; Shi & Malik, 2000; Bianchi et al., 2020), aiming to minimize the number of edges (or similar metric) in a cut, may fall short in capturing the true community structure if the cut’s edge count doesn’t significantly deviate from random graph expectations. This approach originated from the Fiedler vector, which yields a graph cut with a minimal number of edges over all possible cuts (Fiedler, 1973; Newman, 2006a). Similarity-based techniques, reliant on pairwise node similarities, group nodes with shared characteristics. These can be computationally intensive and susceptible to noise, yielding suboptimal results, especially in sparse or noisy data scenarios. Modularity-based methods rely on a statistical approach, measuring the disparity in edge density between a graph and a random graph with the same degree sequence. These modularity maximization methods exhibit a resolution limit (Fortunato & Barthélemy, 2007), as they may inadvertently lead to the neglect of smaller community structures within the graph. Each clustering approach, whether cut-based, similarity-based, or modularity-based, bears its own set of limitations that necessitates careful consideration based on the characteristics of the underlying graph data.
Moreover, graph reduction techniques such as coarsening, summarization, or condensation can also be utilized to facilitate the clustering task (Dhillon et al., 2007; Kumar et al., 2023; Loukas & Vandergheynst, 2018; Loukas, 2019). In the context of graph coarsening, the objective is to learn a reduced graph by merging similar nodes into supernodes. While this coarsening process is traditionally employed for graph reduction, it can be extended to clustering by reducing the original graph such that each class corresponds to a supernode. However, relying solely on coarsening may lead to suboptimal performance, particularly when the order of the reduction from the original graph size to the number of classes is large and results in a significant loss of information.
In this study, we introduce an optimization-based framework designed to enhance clustering by leveraging both the adjacency and the node features of the graph. Our proposed framework strategically incorporates coarsening and modularity maximization, refining partitioning outcomes and bolstering the effectiveness of the clustering process. The algorithm minimizes a nuanced loss function, Q-MAGC objective, encompassing a $\log\det$ term, smoothness, and modularity components, to ensure efficient clustering. Formulated as a multi-block non-convex optimization problem, our approach is adeptly addressed through a block majorization-minimization technique, wherein variables are updated individually while keeping others fixed. The resulting algorithm demonstrates convergence, showcasing its efficacy in efficiently solving the proposed optimization problem.
@@
We introduce two additional algorithms, Q-VGAE and Q-GMM-VGAE, incorporating variational graph auto-encoders to further enhance clustering accuracy. Through comprehensive experiments, we demonstrate the effectiveness of our proposed algorithms, surpassing the performance of state-of-the-art methods on synthetic and real-world benchmark datasets. Our approach showcases notable improvements in clustering performance, solidifying the robustness and superiority of our proposed methods.
@@
Notations. Let $G=\{V,E,A,X\}$ be a graph with node set $V=\{v_{1},v_{2},...,v_{p}\}\ (|V|=p)$, edge set $E\subset V\times V\}(|E|=e)$, weight (adjacency) matrix $A$ and node feature matrix $X\in\mathbb{R}^{p\times n}$. Also, let $\textbf{d}=A\cdot\textbf{1}_{p}\in\mathbb{Z}_{+}^{p}$ be the degree vector, where $\textbf{1}_{p}$ is the vector of size $p$ having all entries 1. Then, the graph Laplacian is $\Theta=\textrm{diag}(\textbf{d})-A$ and the set of all valid Laplacian matrices is defined as: $S_{\Theta}=\{\Theta\in\mathbb{R}^{p\times p}|\Theta_{ij}=\Theta{ji}\leq 0\ \text{for}\ i\neq j,\Theta_{ii}=\sum_{j=1}^{p}\Theta_{ij}\}$.
@@ 2 Related Works
Graph Clustering via Coarsening. Graph coarsening can be extended to graph clustering by reducing the size of the coarsened graph to the number of classes ($k$). However, in most graphs, the number of classes is very small, and reducing the graph to this extent may lead to poor clustering quality. DiffPool (Ying et al., 2018) learns soft cluster assignments at each layer of the GNN and optimizes two additional losses, an entropy to penalize the soft assignments and a link prediction based loss. Next, SAGPool (Lee et al., 2019) calculates attention scores and node embeddings to determine the nodes that need to be preserved or removed. Top-k (Gao & Ji, 2019) also works by sparsifying the graph with the learned weights. MinCutPool (Bianchi et al., 2020) formulates a differentiable relaxation of spectral clustering via pooling. However, Tsitsulin et al. (2023) show that MinCutPool’s orthogonal regularization dominates over the clustering objective and the objective is not optimized. Some disadvantages of these methods are instability and computational complexity in the case of SAGPool and DiffPool and convergence in MinCutPool. To address these challenges, we need a better loss function to perform the clustering task effectively.
@@
Deep Graph Clustering. Previous literature can be classified based on contrastive and non-contrastive methods. On the non-contrastive side, Pan et al. (2018) proposed ARGA and ARVGA, enforcing the latent representations to align to a prior using adversarial learning. By utilizing an attention-based graph encoder and a clustering alignment loss, Wang et al. (2019) propose DAEGC. Liu et al. (2022) design the DCRN model to alleviate representation collapse by a propagation regularization term minimizing the Jensen Shannon Divergence (JSD) between the latent and its product with normalized $A$. Contrastive methods include AGE (Cui et al., 2020) which builds a training set by adaptively selecting node pairs that are highly similar or dissimilar after filtering out high-frequency noises using Laplacian smoothing. Zhao et al. (2021) propose GDCL to correct the sampling bias by choosing negative samples based on the clustering label.
@@
VGAEs (Kipf & Welling, 2016a) are an increasingly popular class of GNNs that leverage variational inference (Kingma & Welling, 2014) for learning latent graph representations in unsupervised settings. They reconstruct the adjacency matrix after passing the graph through an encoder-decoder architecture. Many attempts have been made to use VGAEs with k-means on latent embeddings, but it has been unsuitable for clustering. This is primarily because embedded manifolds obtained from VGAEs are curved and must be flattened before any clustering algorithms using Euclidean distance are applied. Refer to Appendix Section L for a detailed explanation. VGAEs only use a single Gaussian prior for the latent space, whereas clustering requires the integration of meta-priors. Additionally, the inner-product decoder fails to capture locality and cluster information in the formed edges. Several clustering-oriented variants of VGAEs (Mrabah et al., 2022; Hui et al., 2020) have been developed that overcome most of these challenges. GMM-VGAE (Hui et al., 2020) partitions the latent space using a Gaussian Mixture Model and assigns a separate prior for each cluster to better model complex data distributions. Despite the improvement in performance, it’s inner-product decoder cannot capture locality information.
@@
Modularity Maximization. Various heuristic algorithms have been established that solve the NP-hard problem of modularity maximization including sampling, simulated annealing (Guimerà & Nunes Amaral, 2005; Guimerà & Amaral, 2005), and greedy algorithms (Louvain/Leiden) (Newman, 2004; Blondel et al., 2008). These algorithms require intensive compute and don’t use node features. Modularity maximization using GNNs has also garnered attention recently. DMoN (Tsitsulin et al., 2023) optimizes only for modularity with a collapse regularization to prevent the trivial solution, but offers no theoretical guarantees about convergence. Modularity-Aware GAEs and VGAEs (Salha-Galvan et al., 2022) use a prior membership matrix using Louvain algorithm and optimize for modularity using an RBF kernel as a proxy for same-community assignment. DGCLUSTER (Bhowmick et al., 2023) is a semi-supervised method that makes use of either a subset of labels or pairwise memberships as Auxiliary Information coupled with modularity.
@@ 3.1 Graph Coarsening
Graph coarsening is a graph dimensionality reduction technique used in large-scale machine learning to construct a smaller or coarsened graph $G_{c}$ from the original graph $G=\{V,E,A,X\}$ while preserving properties of the original graph $G$. Graph coarsening aims to learn a mapping matrix $C\in\mathbb{R}_{+}^{p\times k}$, where $p$ is the number of nodes in the original graph and $k$ is the number of nodes in the coarsened graph. Each non-zero entry of the mapping matrix $C$, i.e., $C_{ij}$, indicates that the $i$-th node of $G$ is mapped to the $j$-th supernode. Moreover, for a balanced mapping, the mapping matrix $C$ belongs to the following set (Kumar et al., 2023; Loukas & Vandergheynst, 2018; Loukas, 2019):
@@
For $C\in\mathcal{C}$, the relationship between the original graph Laplacian $\Theta$, the coarsened graph Laplacian $\Theta_{c}$, and the mapping matrix $C$ is given by $\Theta_{c}=C^{T}\Theta C$.
@@ 3.2 Spectral Modularity Maximization
Spectral Clustering is the most direct approach to graph clustering, where we minimize the volume of inter-cluster edges. Modularity, introduced in Newman (2006b), is the difference between the number of edges within a cluster $\mathbf{C_{i}}$ and the expected number of such edges in a random graph with an identical degree sequence. It is mathematically defined as:
@@
$B$ $=A-\frac{\textbf{dd}^{T}}{2e},\qquad\textbf{d}=A\cdot\mathbbm{1},\qquad\mathcal{Q}=\frac{1}{2e}Tr(C^{T}BC)$ (3)
$B$ is symmetric and is defined such that its row-sums and column-sums are zero, thereby making $\mathbf{1}$ one of its eigenvectors and 00 the corresponding eigenvalue. These spectral properties of the modularity matrix are also observed in the Laplacian, as noted in Newman (2006b), which is a crucial element in spectral clustering. Modularity is maximized when $u_{1}^{T}s$ is maximized, where $u$ are the eigenvectors of $B$ and $s$ is the community assignment vector, i.e., placing the majority of the summation in $Q$ on the first (and largest) eigenvalue of $B$. Moreover, modularity is closely associated with community detection. These special spectral properties make $B$ an ideal choice for graph clustering. While modularity maximization has been extensively studied, heuristic algorithms for it are computationally intensive, such as the Newman-Girvan algorithm with $\mathcal{O}(p^{3})$ time complexity. The Louvain/Leiden algorithms (Newman, 2004; Blondel et al., 2008) improve on this.
@@ 3.3 Motivation and Proposed Problem Formulation
Current graph clustering methods often struggle to accurately capture both intra-cluster and inter-cluster relationships, achieve computational efficiency, and identify smaller communities, thereby limiting their effectiveness. These methods face significant challenges including instability and the lack of convergence guarantees, especially when employing coarsening techniques. Modularity maximization, despite extensive study, relies on computationally intensive heuristics and lacks theoretical convergence guarantees. To address these issues, we propose an optimization-driven framework that strategically integrates coarsening and modularity. By incorporating both adjacency and node features, our approach aims to robustly capture intra-cluster and inter-cluster dynamics. This framework ensures stability, guarantees convergence, and consistently delivers superior clustering results with enhanced computational efficiency compared to existing methods. Given original graph $G(V,E,A,X)$, the proposed generic optimization formulation is:
@@
subject $\text{subject to}\ C$ $\in\mathcal{S_{C}}=\{C\in\mathbb{R}^{p\times k}|C\geq 0,\ \left\|[C^{T}]_{i}\right\|_{2}^{2}\leq 1\}\;\ \forall\ i=1,2\ldots p$ (4)
Here, $C\in\mathbb{R}_{+}^{p\times k}$ represents the clustering matrix to be learned, where each non-zero entry $C_{ij}$ indicates that the $i$-th node is assigned to the $j$-th cluster. The term $f(C,X,\Theta)$ denotes the graph coarsening objective, managing inter-cluster edges. The function $g(C,A)$ represents the modularity objective, improving clustering performance. Additionally, $h(C,\Theta)$ is a regularization term enforcing desirable properties in the clustering matrix, as defined in Equation 1. The overarching goal of the optimization problem 3.3 is to choose $f(C,X,\Theta)$, $g(C,A)$, and $h(C,\Theta)$ such that nodes are optimally clustered and inter-cluster connectivity is maintained. In the next section, we will develop the clustering algorithm, leveraging the coarsening objective and modularity to enhance clustering performance.
@@ 4 Proposed Method
Given a graph $G=\{V,E,A,X\}$, considering $f(C,X_{C},\Theta)=\text{tr}(X_{C}^{T}C^{T}\Theta CX_{C})$, $g(C,A)=-\frac{\beta}{2e}\textrm{tr}(C^{T}BC)$ (where $B=A-\frac{\textbf{dd}^{T}}{2e},\quad\textbf{d}=A\cdot\mathbbm{1}$), and $h(C,\Theta)=-\gamma\log\det(C^{T}\Theta C+J)$, to obtain clustering matrix $C$ we formulate the following optimization problem :
@@
subject $\text{subject to}\ C\in\mathcal{S_{C}}$ =$=\{C\in\mathbb{R}^{p\times l}\big{|}\ \left\|C^{T}_{i}\right\|_{2}^{2}\leq 1\}\;\forall i,X=CX_{C}\;\;\text{where, }J=\frac{1}{k}\textbf{1}_{k\times k}$ (5)
The term $\textrm{tr}(X_{C}^{T}C^{T}\Theta CX_{C})$ represents the smoothness or Dirichlet energy of the coarsened graph while $C^{T}\Theta C$ is the Laplacian matrix of the coarsened graph. Minimizing smoothness ensures that the clusters or supernodes with similar features are linked with stronger weights. The term $\textrm{tr}(C^{T}BC)$ corresponds to the graph’s modularity, enhancing the quality of the clusters formed. The term $-\log\det(C^{T}\Theta C+J)$ is crucial for maintaining inter-cluster edges in the coarsened graph. For a connected graph matrix with $k$ super-nodes or clusters, the rank of $C^{T}\Theta C$ is $k-1$. Adding $J$ to $C^{T}\Theta C$ makes $C^{T}\Theta C+J$ a full-rank matrix without altering the row and column space of $C^{T}\Theta C$ (Kumar et al., 2020).
Problem 4 is a multi-block non-convex optimization problem solved using the Block Successive Upper-bound Minimization (BSUM) framework. All terms except modularity are convex in nature, which we prove in Appendix A. We iteratively update $C$ and $X_{C}$ alternately while keeping the other constant. This process continues until convergence or the stopping criteria are met. Since the constraint $X=CX_{C}$ is hard and difficult to enforce, we relax it by adding the term $\frac{\alpha}{2}\|X-CX_{C}\|_{F}^{2}$ to the objective. This term ensures that each node is assigned to a cluster, leaving no node unassigned. Additionally, when needed, we can add an optional sparsity regularization term $\lambda\|C^{T}\|_{1,2}^{2}$, which can be seen as ensuring that each node is assigned to exactly one cluster, avoiding any overlap in node assignments across clusters.
@@
By using the first-order Taylor series approximation, a majorised function for $f(C)$ at $C^{t}$ ($C$ after $t$ iterations) can be obtained as:
@@
To check whether the proposed objective (Equation 4) results in consistent clustering, we need to assume a random graph. A graph $G(V,E)$ is generated by a DC-SBM with $p$ nodes which belong to $k$ classes. Following the setup in Zhao et al. (2012), each node $v_{i}$ is associated with a label-degree pair $(y_{i},t_{i})$ drawn from a discrete joint distribution $\Pi_{k\times m}$. Here, $t_{i}$ takes values $0\leq x_{1}\leq\dots\leq x_{m}$ with $\mathbb{E}[t_{i}]=1$, and $y_{i}\in[k]$.
@@
We define $O(e)\in\mathbb{Z}_{\geq 0}^{k\times k}$ as the inter-cluster edge count matrix for a given cluster assignment $e\in[k]^{p}$. Here, $O_{ql}(e)=\sum_{ij}A_{ij}\mathbbm{1}\{e_{i}=q,e_{j}=l\}$ denotes the number of edges between clusters $q$ and $l$. Also, We represent the class frequency distribution as $\pi\in[0,1]^{k}$, where $\pi_{i}$ indicates the fraction of nodes assigned to cluster $i$.
@@
(Theorem 4.1 from Zhao et al. (2012)) For any $\mathcal{L}(e)=F(\frac{O(e)}{\mu_{p}},f(e))$, if $F$ is uniquely maximized by $S=\mathbb{D}$ and $\pi,P,F$ satisfy the regularity conditions, then $\mathcal{L}$ is strongly consistent under DC-SBMs if $\frac{\lambda_{p}}{logp}\rightarrow\infty$ and weakly consistent if $\lambda_{p}\rightarrow\infty$:
@@
It is important to note that the paper focuses on maximizing an objective function. Therefore, we consider the negative of our loss function, $-\mathcal{L}_{MAGC}$. Additionally, we only consider the objective function defined in equation 4, as regularizers can be adjusted for different downstream tasks.
@@
We also demonstrate experimental consistency of the optimization objective equation 4 in Appendix I, resulting in complete recovery of the labels. This consistency extends to the objective with additional regularization terms, as detailed therein. We want to emphasize that the consistency analysis for Equation 4 holds regardless of how the objective is optimized: whether by integrating our loss into Graph Neural Networks (GNNs) or by employing Block Majorization-Minimization techniques.